# Dephasing and error dynamics affecting a singlet-triplet qubit during coherent spin shuttling

### v1

@@ IV.1 Device details
The device was fabricated by Intel Corporation and is based on a Si/Si0.7Ge0.3 heterostructure Neyens2024 , where the quantum well is composed of isotopically-purified 28Si (800 ppm 29Si). Gate-defined QDs are formed by molding the local electrostatic potential with voltages applied to the surrounding gates. A reservoir of electrons is hosted under the accumulation gate AC0 from which the QDs are populated. Changes in charge occupation are sensed by a single electron transistor (SET) located at a dedicated QD on the top half of the device opposite the qubit QDs and isolated by a center screening gate. The dot pitch is $\sim$120 nm and the well width is $\sim$5 nm.
@@ IV.2 Experimental setup
We used a closed-loop dry cryostat Bluefors LD400 dilution refrigerator with a base temperature of 7 mK. The electron temperature was approximately 250 mK. We used a heterojunction bipolar transistor (HBT) operating at -1.056 V emitter bias on the readout SET line to perform cryogenic preamplification Curry2019 . DC voltage to the gates was supplied by a QDevil QDAC II and fast AC pulses were generated by a Keysight M3201 AWG with a 500 MHz sampling rate. DC and fast voltage pulses are combined through a cryogenic RC bias tee at the PCB. Each individual spin state measurement is acquired over an average of 100 shots. We used one axis of a superconducting vector magnet aligned to the [110] crystal direction in the plane of the device, perpendicular to the direction of shuttling.

### v2

@@ IV.1 Device details
The device was fabricated by Intel Corporation and is based on a Si/Si0.7Ge0.3 heterostructure Neyens2024 , where the quantum well is composed of isotopically-purified 28Si (800 ppm 29Si). Gate-defined QDs are formed by molding the local electrostatic potential with voltages applied to the surrounding gates. A reservoir of electrons is hosted under the accumulation gate AC0 from which the QDs are populated. Changes in charge occupation are sensed by a single electron transistor (SET) located at a dedicated QD on the top half of the device opposite the qubit QDs and isolated by a center screening gate. The dot pitch is $\sim$120 nm, the Si well thickness is $\sim$5 nm, and the SiGe barrier thickness is $\sim$50 nm.
@@ IV.2 Experimental setup
We use a closed-loop dry cryostat Bluefors LD400 dilution refrigerator with a base temperature of 7 mK. The electron temperature is approximately 250 mK. We use a heterojunction bipolar transistor (HBT) operating at -1.056 V emitter bias on the readout SET line to perform cryogenic preamplification Curry2019 . DC voltage to the gates is supplied by a QDevil QDAC II and fast AC pulses are generated by a Keysight M3201 AWG with a 500 MHz sampling rate. DC and fast voltage pulses are combined through a cryogenic RC bias tee at the PCB. Each individual spin state measurement is acquired over an average of 100 shots. We use one axis of a superconducting vector magnet aligned to the [110] crystal direction in the plane of the device, perpendicular to the direction of shuttling.